OpenStudy (pokemon23):

If the square of a certain number added to its square is 30.Find the number.

OpenStudy (pokemon23):

\[x+x^2=30\]

OpenStudy (anonymous):

30^2

OpenStudy (anonymous):

x^2+x-30=0 (x+6)(x-5)=0 x-5=0

OpenStudy (anonymous):

and x+6=0

OpenStudy (pokemon23):

isn't it supposed to equal 30?

OpenStudy (anonymous):

the number when added to its square is supposed to be equal to 30.

OpenStudy (angela210793):

I thought it was x^2+x^2=30 ....O.o

OpenStudy (pokemon23):

so shouldn't it be x^2+x^2=30 because it says added to its square is 30 which mean equals 30?

OpenStudy (pokemon23):

@dumbcow

OpenStudy (pokemon23):

@Directrix

OpenStudy (dumbcow):

"added to its square" i guess they mean the square of the number squared --> (x^2)^2 = x^4 x^4 + x^2 = 30 x^4 + x^2 -30 = 0 (x^2 -5)(x^2 +6) = 0 ...

OpenStudy (angela210793):

isn't the square of a no. x^2???? O.o

OpenStudy (dumbcow):

haha yes, but "it" is referring to the square of a number

OpenStudy (pokemon23):

I see...

OpenStudy (anonymous):

oh

OpenStudy (pokemon23):

I learn something today..

OpenStudy (angela210793):

Hmmmmm.....i still don't get it -_-

OpenStudy (dumbcow):

we need translators for some of the word problems out there....its not english class gosh textbook writers c'mon

OpenStudy (anonymous):

lets assume x is already some square and its added it its square.

OpenStudy (angela210793):

why do u have to assume that????????? O.o

OpenStudy (pokemon23):

blame my teacher she write these silly grammar problems

OpenStudy (anonymous):

Ok, lets not assume, lets take it for a fact.

OpenStudy (angela210793):

I got it ...I got it ^_^

OpenStudy (pokemon23):

Makes me chuckle

OpenStudy (anonymous):

but still the solution doesnt make sense lol

OpenStudy (dumbcow):

2 of the answers are imaginary, so yes those don;t make sense but the other 2 are +-sqrt5

OpenStudy (anonymous):

yeah ^

OpenStudy (angela210793):

yea..u can subtitude x^2 with ''t'' and then solve it as a second grade equation

OpenStudy (pokemon23):

I will ask my silly teacher about solving this problem

OpenStudy (anonymous):

ask her is +-sqrt5 is correct.